# Preference-Guided Reinforcement Learning for Efficient Exploration

### v1

@@ IV-B Two-step Policy Optimization Framework
This section presents a two-step policy optimization framework to provide exploration aid and accelerate learning for the agent. Formally, the policy $\pi_{\theta}$ is updated by the following process:
@@
LOPE performs one-step policy improvement by optimizing the standard trust-region-based objective function [51]. This can be expressed as follows:
@@
[…]
Here, $d_{k}$ is the discounted state visitation distribution corresponding to $\pi_{k}$, and $A_{k}(s,a)$ is the advantage of $\pi_{k}$ in the state-action pair $(s,a)$. The trust-region-based update in Eq. (7) is to find a new policy $\pi_{k+\frac{1}{2}}$ that maximizes the RL objective while ensuring this new policy stays within the trust region around $\pi_{k}$. From an implementation perspective, off-the-shelf algorithms can solve this optimization problem, such as TRPO [51] or PPO [52].
@@
Many RL approaches achieve efficient online policy optimization in dense reward settings. Nonetheless, they may often fail due to the sparsity of reward functions. It is a forlorn hope to achieve significant performance improvement with sparse-reward feedback in the initial training stage. We propose to overcome this difficulty of exploration by using preference as guidance and adjusting policy optimization directly. The preference guidance step is given as follows:
@@
[…]
This preference guidance step is derived from Eq. (6). We solve Eq. (8) by introducing a novel policy gradient. Instead of directly computing the gradient update direction to maximize Eq. (8) as [48], we seek to solve an equivalent optimization problem, which is obtained by expanding the objective function in Eq. (8). This divergence minimization problem can be reduced into a policy-gradient algorithm with shaped rewards computed from $\mathcal{P}$.
@@
$r_{g}(s,a)=\underset{\begin{subarray}{c}\tau\in\mathcal{P}_{s,a}\end{subarray}}{\mathbb{E}}\left[\mathrm{dist}(\tau,\mathcal{P})\right].$ (9)
Here, $\mathrm{dist}(\tau,\mathcal{P})$ is considered the distance between $\tau$ and $\mathcal{P}$, i.e. $\mathrm{dist}(\tau,\mathcal{P})=\mathbb{E}_{\upsilon\sim\mathcal{P}}\left[d(\tau,\upsilon)\right]$, and $\mathcal{P}_{s,a}=\{\tau\in\mathcal{P}|(s,a)\in\tau\}$. Then, Eq. (8) can be expanded as follows:
@@
We provide the proof in Appendix A. Compared to offline demonstrations as guidance [28], our approach considers the human-in-the-loop setting where the agent obtains preference-labeled trajectories by interacting with the expert annotator. Therefore, although sparse rewards are unavailable for the agent in the initial training phase, LOPE can orient policy optimization by capturing more task-related information from human feedback. On the other hand, if the support sets of $\pi$ and $\pi_{b}$ are far apart or do not overlap at all, then the value of KL divergence is meaningless. In this case, the method of [28] does not work, but we provide a feasible approach to overcome the gradient vanishing problem. The entire procedure of LOPE is summarized in Algorithm 1. Hence, compared with the result of Lemma 1 in the literature [28], our result of Lemma 2 has better generalizability because of the continuity of the MMD distance.
To solve the optimization problem in Eq. (10), all that remains is to replace the expectation with sample averages, and we use the preference guidance rewards to estimate the expected returns. In this manner, existing algorithms, such as PPO or TRPO, can now easily solve the optimization problem in Eq. (8). The entire procedure of LOPE is summarized in Algorithm 1.
@@
LOPE maintains a buffer of preferred trajectories and updates this buffer with current trajectories in each iteration. Therefore, the human annotator must compare all current trajectories with the preferred trajectories in $\mathcal{P}$ in pairs. The trajectories are ranked by human feedback, and the best $h$ trajectories (segments) can be stored in $\mathcal{P}$. In practice, we set $h=8$ in all experiments. To achieve the $n$-wise comparisons, each current trajectory should be compared pair-wisely with all preference trajectories in $\mathcal{P}$. When the agent samples $N$ new trajectories in each epoch, the computational complexity of $n$-wise comparisons is $Nh$. As the training progresses, the number of trajectories used to update $\mathcal{P}$ decreases to reduce the experts’ workload. The decreased number is environmentally dependent and relies on the agent’s training progress.
@@ V Theoretical Analysis
In this section, we present the theoretical foundation of LOPE and derive a worst-case lower bound on the policy performance improvement for LOPE. We analyze the policy improvement step and preference guidance step separately. First, similar to [51], we introduce the following local approximation to $J(\cdot)$ as follows:
$L_{\pi}(\tilde{\pi})=J(\pi)+\sum_{s}\rho_{\pi}(s)\sum_{a}\tilde{\pi}(a|s)A_{\pi}(s,a),$ (11)
where $\rho_{\pi}$ is the discounted state visitation distribution of $\pi$, and $A_{\pi}$ is the advantage function of $\pi$. We then make the following assumption about the behavior policy $\pi_{b}$ implied by the preferred trajectories in set $\mathcal{P}$.
@@
where $\rho_{k}$ is the discounted state visitation distribution of $\pi_{k}$, and $L_{k}$ is the locally linear approximation to $J(\cdot)$ near $\pi_{k}$.
Figure 2: The four environments for evaluating the LOPE’s performance: (a) Grid World; (b) SparseHalfCheetah; (c) SparseHopper; (d) Ant Maze.
Eq. (13) implies that the agent will produce a higher advantage by performing actions according to the demonstrations in $\mathcal{P}$. This is a relatively reasonable assumption since the demonstration preferred by the annotator is likely to achieve better performance than a policy that is not fully trained. It is worth mentioning that a similar assumption is adopted by [53] and [28]. However, different from these methods, $\pi_{b}$ in our study is not the final optimal policy, and the expected return of $\pi_{b}$ may gradually increase as an agent stores more preferred trajectories with high returns.
We begin with the analysis of the policy improvement step, and the following conclusion and its analysis are standard in the related literature [54].
Proposition 1 (Performance improvement bound of the PI step [54]).
@@
Lemma 2 provides a quantitative description of performance improvement in the preference-guided step. The human annotator selects preferred trajectories and stores them in $\mathcal{P}$ during training. Hence, it is reasonable to assume that $\pi_{k+\frac{1}{2}}$ in the PG step satisfies Assumption 1 in the preference-guided step. Based on these preferred trajectories, the agent obtains a higher worst-case lower bound of performance improvement. Specifically, according to Lemma 2, for $\pi=\pi_{k+\frac{1}{2}}$, Eq. (15) indicates that the new lower bound increases by $\Delta/(1-\gamma)$ than in cases where preferred trajectories are not provided.
We derive a new worst-case lower bound of LOPE’s performance improvement by combining Proposition 1 and Lemma 2.
@@
We provide the proof in Appendix B. An intuitive explanation of this theorem is given in the following remark.
Remark 3.
When the behavior policy $\pi_{b}$ performs much better than the current policy as described in Assumption 1, compared to the original worst-case lower bound obtained by the TRPO algorithm, the preference-guided step results in an additional performance improvement of $\Delta/(1-\gamma)$ size. Preference-labeled trajectories provide a more effective and accurate policy optimization direction with the agent while decreasing sampling complexity for learning. Thus, LOPE can achieve faster learning.
@@ VI Experimental Setups
We evaluate LOPE on an enormous grid-world task and several continuous control tasks based on the MuJoCo physical engine [55]. The hyper-parameters used by the PyTorch code are introduced in Appendix C. In our experiments, sparse rewards are only produced by the environments, and the agent has access to both sparse rewards and human feedback.
@@ VI-A Environments
Key-Door-Treasure domain. As shown in Figure 2a, the Key-Door-Treasure domain is a grid-world environment similar to that defined by [7] with minor modifications. The agent should pick up the key (K) to open the door (D) and finally collect the treasure (T). The agent can only obtain a reward of 200 when reaching the treasure, and it cannot receive any rewards in other cases. The size of this grid-world environment is $26\times 36$.
Locomotion control tasks. To explore the limitation of LOPE, we modified 2 MuJoCo tasks, HalfCheetah and Hopper, and obtained two new robots: SparseHalfCheetah and SparseHopper presented in Figs. 2b and 2c. Each of the new robots yielded a sparse reward only when the center-of-mass of the corresponding robot was beyond a certain threshold distance against the initial position, and the agent cannot be rewarded in other cases.
Ant Maze. To further demonstrate the feasibility of LOPE, we evaluated LOPE on an Ant-Maze task designed as a benchmark for RL by [56]. Hard exploration manifests two aspects: Locomotion and Maze. More specifically, the agent must first learn to walk and then reach the target point. The Ant is only rewarded for reaching the specified position in a maze shown in Fig. 2d. The observation space of this task is naturally decomposed into two parts: the agent’s joint angle information and task-specific attributes. The task-specific attributes include sensor readings and the positions of walls and goals.
@@ VI-B Baseline Methods
For evaluation, we compare LOPE with PEBBLE [13], the current state-of-the-art PbRL method. Additionally, our method aims to overcome exploration difficulties; therefore, NoisyNet [57] and PPO+EXP [58] are adopted as baseline methods. Meanwhile, SIL [29] and GASIL [59] are viewed as two baselines to demonstrate the LOPE’s effectiveness in exploiting previous good experiences. Lastly, we use the standard PPO [52] as a baseline method.
PEBBLE is the current state-of-the-art PbRL method. PEBBLE uses an entropy-based sampling method to improve sample and exploration efficiency. PEBBLE adopts the setting where the agent can only access expert preferences $y$ for segments $\sigma^{0}$ and $\sigma^{1}$. Specifically, $y\in\{0,0.5,1\}$ is sampled from a discrete distribution to indicate which segment the human annotator prefers. Then, the agent records the judgment as a triple $(\sigma^{0},\sigma^{1},y)$ in a dataset $\mathcal{D}$. Each time the dataset is updated, the reward model is trained to remain consistent with the preference data.
Compared with PEBBLE, the primary differences in our approach are (1) an end-to-end PbRL formulation, (2) the subtle combination of online optimization and preference guidance, and (3) the preferred set update method with $n$-wise comparison.

### v2

@@ IV-B Two-step Policy Optimization Framework
This section presents a two-step policy optimization framework that provides exploration aid and accelerates learning for the agent. Formally, the policy $\pi_{\theta}$ is updated by the following process:
@@
LOPE performs one-step policy improvement by optimizing the standard trust-region-based objective function [schulman2015trust]. This can be expressed as follows:
@@
$\text{s.t.}\ D_{\rm KL}(\pi_{\theta},\pi_{k})\leq\delta.$
Here, $d_{k}$ is the discounted state visitation distribution corresponding to $\pi_{k}$, and $A_{k}(s,a)$ is the advantage of $\pi_{k}$ in the state-action pair $(s,a)$. The trust-region-based update in Eq. (7) is to find a new policy $\pi_{k+\frac{1}{2}}$ that maximizes the RL objective while ensuring this new policy stays within the trust region around $\pi_{k}$. From an implementation perspective, off-the-shelf algorithms can solve this optimization problem, such as TRPO [schulman2015trust] or PPO [schulman2017proximal].
@@
Many RL approaches achieve efficient online policy optimization in dense reward settings. Nonetheless, they may often fail due to the sparsity of reward functions. It is a forlorn hope to achieve significant performance improvement with sparse-reward feedback in the initial training stage. We propose overcoming this difficulty in exploration by using preference as guidance and adjusting policy optimization directly. The preference guidance step is given as follows:
@@
$\text{s.t.}\ D_{\rm KL}(\pi_{\theta},\pi_{k+\frac{1}{2}})\leq\delta.$
This preference guidance step is derived from Eq. (6). We solve Eq. (8) by introducing a novel policy gradient. Instead of directly computing the gradient update direction to maximize Eq. (8) as [masood2019diversity], we seek to solve an equivalent optimization problem, which is obtained by expanding the objective function in Eq. (8). This divergence minimization problem can be reduced to a policy-gradient algorithm with shaped rewards computed from $\mathcal{P}$.
Noting that the equation of MMD in Eq. (5) can be rewritten as:
$d(\tau,\upsilon)=$ $\underset{s_{\tau}\sim\rho_{\tau}}{\mathbb{E}}\Big[\underset{s_{\tau}^{\prime}\sim\rho_{\tau}}{\mathbb{E}}\left[k\left(s_{\tau},s_{\tau}^{\prime}\right)\right]-2\underset{s_{\upsilon}\sim\rho_{\upsilon}}{\mathbb{E}}\left[k(s_{\tau},s_{\upsilon})\right]$ (9)
$+\underset{s_{\upsilon},s_{\upsilon}^{\prime}\sim\rho_{\upsilon}}{\mathbb{E}}\left[k(s_{\upsilon},s_{\upsilon}^{\prime})\right]\Big].$
Hence, the definition of $d(\cdot,\cdot)$ can be regarded as taking expectations of the distance from the state $s_{\tau}$ to the trajectory $\upsilon$:
$d(\tau,\upsilon)=$ $\underset{s_{\tau}\sim\rho_{\tau}}{\mathbb{E}}\left[\operatorname{dist}(s_{\tau},\upsilon)\right],$ (10)
where $\operatorname{dist}(s_{\tau},\upsilon)$ is computed as:
$\operatorname{dist}(s_{\tau},\upsilon)=$ $\underset{s_{\tau}^{\prime}\sim\rho_{\tau}}{\mathbb{E}}\left[k\left(s_{\tau},s_{\tau}^{\prime}\right)\right]-2\underset{s_{\upsilon}\sim\rho_{\upsilon}}{\mathbb{E}}\left[k(s_{\tau},s_{\upsilon})\right]$ (11)
$+\underset{s_{\upsilon},s_{\upsilon}^{\prime}\sim\rho_{\upsilon}}{\mathbb{E}}\left[k(s_{\upsilon},s_{\upsilon}^{\prime})\right]$
@@
$r_{g}(s,a)=\underset{\begin{subarray}{c}\upsilon\in\mathcal{P}\end{subarray}}{\mathbb{E}}\left[\mathrm{dist}(s,\upsilon)\right].$ (12)
$\operatorname{dist}(\cdot,\cdot)$ is defined in Eq. (11). Then, Eq. (8) can be expanded as follows:
@@
We provide the proof in Appendix A. Compared to offline demonstrations as guidance [rengarajanreinforcement], our approach considers the human-in-the-loop setting where the agent obtains preference-labeled trajectories by interacting with the expert annotator. Therefore, although sparse rewards are unavailable to the agent during the initial training phase, LOPE can guide policy optimization by capturing more task-related information from human feedback. On the other hand, if the support sets of $\pi$ and $\pi_{b}$ are far apart or do not overlap at all, then the value of the KL divergence is meaningless. In this case, the method of [rengarajanreinforcement] does not work, but we provide a feasible approach to overcome the gradient vanishing problem. The entire procedure of LOPE is summarized in Algorithm 1. Hence, compared with the result of Lemma 1 in the literature [rengarajanreinforcement], our result of Lemma 2 has better generalizability because of the continuity of the MMD distance.
To solve the optimization problem in Eq. (13), all that remains is to replace the expectation with sample averages, and we use the preference guidance rewards to estimate the expected returns. In this paper, we implement LOPE based on PPO. PPO utilizes a clipped surrogate objective to implicitly constrain policy updates within a trust region, thereby eliminating the need to explicitly set or tune $\delta$. The entire procedure of LOPE is summarized in Algorithm 1.
@@
LOPE maintains a buffer of preferred trajectories and updates this buffer with current trajectories in each iteration. Therefore, the human annotator must compare all current trajectories with the preferred trajectories in $\mathcal{P}$ in pairs. The trajectories are ranked by human feedback, and the best $h$ trajectories (segments) can be stored in $\mathcal{P}$. In practice, we set $h=8$ in all experiments. To achieve the $n$-wise comparisons, each current trajectory should be compared pair-wise with all preference trajectories in $\mathcal{P}$. When the agent samples $N$ new trajectories in each epoch, the computational complexity of $n$-wise comparisons is $Nh$. As the training progresses, the number of trajectories used to update $\mathcal{P}$ decreases, thereby reducing the experts’ workload. The decreased number is environmentally dependent and relies on the agent’s training progress. Furthermore, our method’s experts only need to assess the overall state visitation situation of the agent’s trajectory, without requiring a detailed understanding of the agent’s dynamics. This reduces the requirements for experts.
@@ V Theoretical Analysis
This section analyzes the policy improvement step and preference guidance step separately. We make the following assumption about the behavior policy $\pi_{b}$ implied by the preferred trajectories in set $\mathcal{P}$.
Figure 2: The advantage of the guiding policy $\pi_{b}$ over the current policy.
@@
where $\rho_{k}$ is the discounted state visitation distribution of $\pi_{k}$.
To empirically validate Assumption 1, we evaluate the guided advantage $\Delta=\mathbb{E}_{a\sim\pi_{b}}\left[A_{k}(s,a)\right]$ in the grid-world maze task with randomized goals. As shown in Fig. 2, the advantage of $\pi_{b}$ over $\pi_{k}$ remains consistently positive and above a threshold ($\approx 0.15$), indicating that the guiding policy reliably satisfies the sufficient condition assumed in our theoretical analysis.
Figure 3: The four environments for evaluating the LOPE’s performance: (a) Grid World; (b) AntMaze-Umaze; (c) PointMaze-Medium; (d) PointMaze-Large.
Eq. (14) implies that the agent will produce a higher advantage by performing actions according to the demonstrations in $\mathcal{P}$. This is a relatively reasonable assumption since the demonstration preferred by the annotator is likely to achieve better performance than a policy that is not fully trained. It is worth mentioning that a similar assumption is adopted by [rengarajanreinforcement] and [kang2018policy]. However, our method does not require an exact estimation of $\rho_{\pi}$ during implementation. Instead, we rely on sampling-based approximations using on-policy trajectories. $\pi_{b}$ serves as a heuristic guide, providing direction for exploration based on human preferences. The expected return of $\pi_{b}$ may gradually increase as the agent stores more preferred trajectories with high returns.
We begin with the analysis of the policy improvement step, and the following conclusion and its analysis are standard in the related literature [achiam2017constrained].
Proposition 1 (Performance improvement bound of the PI step [achiam2017constrained]).
@@
Lemma 2 provides a quantitative description of performance improvement in the preference-guided step. The human annotator selects preferred trajectories and stores them in $\mathcal{P}$ during the training process. Hence, it is reasonable to assume that $\pi_{k+\frac{1}{2}}$ in the PG step satisfies Assumption 1 in the preference-guided step. Based on these preferred trajectories, the agent obtains a higher worst-case lower bound of performance improvement. Specifically, according to Lemma 2, for $\pi=\pi_{k+\frac{1}{2}}$, Eq. (16) indicates that the new lower bound increases by $\Delta/(1-\gamma)$ than in cases where preferred trajectories are not provided.
We derive a new worst-case lower bound of LOPE’s performance improvement by combining Proposition 1 and Lemma 2. The following theorem presents an explicit per-iteration lower bound on the policy improvement achieved by LOPE. This provides a concrete and interpretable expression characterizing the extent of policy improvement as a function of preference advantage, KL divergence, and reward estimation error.
@@
We provide the proof in Appendix B. This bound implies that, when the guided advantage $\Delta$ exceeds the regularization terms, LOPE guarantees monotonic policy improvement. Although the theoretical bound includes a strictly positive term, this quantity serves as a conservative regularization to prevent overly aggressive updates. Our empirical findings indicate that LOPE can still converge to near-optimal solutions in practice.
Specifically, when the behavior policy $\pi_{b}$ performs much better than the current policy as described in Assumption 1, compared to the original worst-case lower bound obtained by the TRPO algorithm, the preference-guided step results in an additional performance improvement of $\Delta/(1-\gamma)$ size. Preference-labeled trajectories offer a more effective and accurate policy optimization direction for the agent, while reducing the sampling complexity required for learning. Thus, LOPE can achieve faster learning.
Remark 3 (Comparison with Prior Preference-Based Bounds).
While most existing preference-based RL methods focus on regret-based frameworks [Pacchiano2021DuelingRR, Chen2022HumanintheloopPE] or reward estimation convergence [biyik2022learning], they typically do not provide formal guarantees on policy improvement. In contrast, LOPE introduces a two-step optimization scheme consisting of a preference-guided advantage improvement objective and a KL-regularized policy update, which together yield the following worst-case lower bound:
$J(\pi_{k+1})-J(\pi_{k})\geq\frac{\Delta}{1-\gamma}-\mathcal{O}\left(\frac{\beta A_{\max}+\sqrt{\delta}A_{\max}}{(1-\gamma)^{2}}\right),$
This bound characterizes the explicit impact of preference quality, policy divergence, and reward variance on performance improvement.
To the best of our knowledge, LOPE is among the first PbRL methods to provide such a per-iteration improvement guarantee without requiring a learned reward function. While similar theoretical guarantees have been studied in the learning-from-demonstration (LfD) literature [rengarajanreinforcement, kang2018policy], they are derived under fundamentally different assumptions, typically requiring access to expert trajectories, cost shaping, or supervised behavior cloning. In contrast, LOPE operates purely at the trajectory level, making our theoretical result novel and better aligned with real-world feedback modalities. These contributions help bridge a theoretical gap in preference-based reinforcement learning.
@@ VI Experimental Setups
We evaluate LOPE on an enormous grid-world task and several continuous control tasks based on the MuJoCo physical engine [todorov2012mujoco]. The hyperparameters used by the PyTorch code are introduced in Appendix C. In our experiments, sparse rewards are generated by the environment, and the agent has access to both sparse rewards and human feedback.
@@ VI-A Environments
Key-Door-Treasure domain. As shown in Figure 3a, the Key-Door-Treasure domain is a grid-world environment similar to that defined by [guo2020memory] with minor modifications. The agent should pick up the key (K) to open the door (D) and finally collect the treasure (T). The agent can only obtain a reward of 400 when reaching the treasure, and it cannot receive any rewards in other cases. The size of this grid-world environment is $26\times 36$. The paper uses mazes with fixed and stochastic goals to test the performance of LOPE. The maze environment with stochastic goals selects a random goal in the upper-right or medium rooms at the beginning of each episode.
Continuous maze tasks. To further demonstrate the feasibility of LOPE, we evaluated LOPE on the continuous maze tasks: PointMaze-Medium, PointMaze-Large, and AntMaze-Umaze task, which are shown in Figs 3b-3d. These tasks are designed as a benchmark for RL by [duan2016benchmarking]. Hard exploration manifests two aspects: Locomotion and Maze. More specifically, the agent must first learn to walk and then reach the target point. The agent is only rewarded for reaching the specified position in each maze. The observation space of this task is naturally decomposed into two parts: the agent’s joint angle or position information and task-specific attributes. The task-specific attributes include sensor readings and the positions of walls and goals.
Figure 4: The two locomotion control tasks: (a) SparseHalfCheetah; (b) SparseHopper.
Locomotion control tasks. To explore the limitation of LOPE, we modified 2 MuJoCo tasks, HalfCheetah and Hopper, and obtained two new robots: SparseHalfCheetah and SparseHopper presented in Figs. 4a and 4b. Each of the new robots yielded a sparse reward only when the center of mass of the corresponding robot was beyond a certain threshold distance from its initial position, and the agent could not be rewarded in any other case.
@@ VI-B Baseline Methods
For evaluation, we compare LOPE with PEBBLE [lee2021pebble] and RUNE [liang2022reward], the current state-of-the-art PbRL methods. Additionally, our method aims to overcome exploration difficulties; therefore, NoisyNet [fortunato2017noisy] and PPO+EXP [tang2017exploration] are adopted as baseline methods. We adopt SIL [oh2018self] and GASIL [guo2018generative] as two baselines to demonstrate the LOPE’s effectiveness in exploiting previous good experiences. Meanwhile, SMM [lee2019efficient] solves the exploration problem by proposing a policy-level state marginal matching objective. We use this method as a baseline to demonstrate the effectiveness of the preference-based trajectory-level state marginal matching objective. Lastly, we use the standard PPO [schulman2017proximal] as a baseline method.
PEBBLE uses an entropy-based sampling method to improve sample and exploration efficiency. RUNE is a PbRL-based exploration method that defines an intrinsic reward by computing the difference between learned reward models. For fairness, PEBBLE and RUNE adopt the setting where the agent can only access expert preferences $y$ for complete trajectories $\tau^{0}$ and $\tau^{1}$. Specifically, $y\in\{0,0.5,1\}$ is sampled from a discrete distribution to indicate which segment the human annotator prefers. Then, the agent records the judgment as a triple $(\tau^{0},\tau^{1},y)$ in a dataset $\mathcal{D}$. Each time the dataset is updated, the reward model is trained to remain consistent with the preference data.
Compared with PEBBLE and RUNE, the primary differences in our approach are (1) an end-to-end PbRL formulation, (2) the subtle combination of online optimization and preference guidance, and (3) the preferred set update method with $n$-wise comparison.